The 2017 balance sheet of Kerbers Tennis Shop, Inc., showed long-term debt of $65 million, and the 2018 balance sheet showed long-term debt of $6.7 million. The 2018 income statement showed an interest expense of $230,000. During 2018, the company had a cash flow to creditors of $30,000 and the cash flow to stockholders for the year was $85,000. Suppose you also know that the firms net capital spending for 2018 was $1,500,000, and that the firm reduced its net working capital investment by $95,000. What was the firms 2018 operating cash flow, or OCF? (Enter your answer in dollars, not millions of dollars, e.g., 1,234,567.) Operating cash flow
Answer #1

Cash From From Assets = Cash flow to Shareholders + Cash flow to Creditors = 85,000 + 30,000 = 105,000

Operating Cash flow = Cash Flow From Assets + Net Capital Spending - Decrease in  Net working Capital = 105,000 + 1,500,000 - 95,000 = 1,510,000

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well